two sides of a triangle have the same length. The third side measures 2m less than twice the common length. The perimeter of the triangle is 14m. what are the lengths of all three sides?
Answer by checkley77(12844)

x+x+2x-2=14
4x=14+2
4x=16
x=16/4
x=4 ans. for the 2 equal sides.
Proof:
4+4+2*4-2=14
4+4+8-2=14
14=14